OpenStudy (anonymous):

The mean of a list of 80 numbers is 230. If the four numbers 145, 156, 210, 255 are added, what is the mean of the list of numbers?

OpenStudy (netlopes1):

Call the sum of 80 numbers "S". Thus,\[(A/80)=230\rightarrow A=18400\] we have the following equation:

OpenStudy (netlopes1):

145+156+210+255=766

OpenStudy (netlopes1):

The new list has 84 numbers and the "new" equation (where the mean I called "M") is:

OpenStudy (netlopes1):

\[(A+766)/84=M \rightarrow or \rightarrow (18400+766)/84=M \rightarrow M \approx 228,17\]

OpenStudy (anonymous):

where do the numbers 84 and 18400 come from?